OpenStudy (anonymous):

Factor the expression completely 1 11 ---x + --- 4 4 A. 1/4(x+5/2) B. 3/4(x+2) C. 1/2 ( 1/2x + 11/2) D. 1/4(x + 11/2)

OpenStudy (alexandervonhumboldt2):

which one?

OpenStudy (anonymous):

I am editing it

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

Done! :3

OpenStudy (alexandervonhumboldt2):

so 11/4=1/2*11/2 1/4x=1/2*1/2x what you think now?

OpenStudy (alexandervonhumboldt2):

both of them share 1/2

OpenStudy (alexandervonhumboldt2):

so write is before the ( )

OpenStudy (funnyguy2):

hell oh

pooja195 (pooja195):

@funnyguy2 thats not necessary to say